OpenStudy (anonymous):

Solve 2(4x + 5) > 7x + 20. {x | x > 2} {x | x < 10} {x | x < 2} {x | x > 10}

OpenStudy (anonymous):

2(4x + 5) > 7x + 20 =8x+10 > 7x + 20 =x+10 > 20 = x > 10

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem ;)